Use the appropriate property of inequality to solve:
7r/12= -14
I tried:
7r/12 x 12 = 12 x (-14)
7r = -2
7r/7=-2/7
r=-2/7
?

Answer by checkley77(12844)

7r/12=-14 cross multiply.
7r=-14*12
7r=-168 divide both sides by 7.
r=-168/7 reduce the fraction.
r=-24 ans.